Suppose that 𝑓 is differentiable and 𝑓′′(π‘Ž) exists. Prove that 𝑓′′(π‘Ž)=limβ„Žβ†’π‘Žπ‘“(π‘Ž+β„Ž)βˆ’2𝑓(π‘Ž)+𝑓(π‘Žβˆ’β„Ž)β„Ž2. Give an example where the above limit exists, but 𝑓′′(π‘Ž) does not exist.


1
Expert's answer
2022-01-06T18:08:04-0500

"f''(x)=\displaystyle \lim_{h\to 0}\frac{f'(x+h)-f'(x)}{h}="


"=\displaystyle \lim_{h\to 0}\frac{\displaystyle \lim_{h_1\to 0}\frac{f'(x+h)-f'(x+h-h_1)}{h_1}-\displaystyle \lim_{h_2\to 0}\frac{f'(x)-f'(x-h_2)}{h_2}}{h}"


letting "h=h_1=h_2", i.e. taking them all to zero at the same rate we have:


"f''(x)=\displaystyle \lim_{h\to 0}\frac{\frac{f'(x+h)-f'(x+h-h)}{h}-\frac{f'(x)-f'(x-h)}{h}}{h}=\displaystyle \lim_{h\to 0}\frac{f(x+h)+f(x-h)-2f(x)}{h^2}"


then:


"f''(a)=\displaystyle \lim_{h\to 0}\frac{f(a+h)+f(a-h)-2f(a)}{h^2}"


example:

for "f'(x)=|x|" :


by L’HΓ΄pital’s rule we have:


"\displaystyle \lim_{h\to 0}\frac{f(x+h)+f(x-h)-2f(x)}{h^2}=\displaystyle \lim_{h\to 0}\frac{f'(x+h)-f'(x)}{2h}="


"=\displaystyle \lim_{h\to 0}\frac{|0+h|-|0-h|}{2h}=0"


That means the limit exists at x = 0, but "f'(x)=|x|" is not differentiable at 0, so "f''(0)"

does not exist.